# PrivacyRestore: Privacy-Preserving Inference in Large Language Models via Privacy Removal and Restoration

### v3

@@ 4.2 Inference Stage
In the inference stage, the client construct a meta vector which is later transmitted to the server along with the input with privacy spans removed, as shown in the lower left panel in Figure 1. These operations are conducted on the client side. Transmitting a single meta vector instead of multiple restoration vectors reduces the communication burden and the risk of data leakage. For instance, adversaries could easily know the number of privacy spans in the user input if multiple restoration vectors were sent.
@@
However, equal weighted aggregation may weaken the influence of critical spans and amplify the effect of irrelevant ones. Therefore, we propose a novel method called Attention-aware Weighted Aggregation (AWA) which estimates a weight for each privacy span, and then take the weighted sum of restoration vectors as the aggregation result. This result is then added with noise for privacy protection and transmitted to the server. Considering the limitation of computing resource, we propose to utilize a lightweight model (e.g., BERT Devlin et al. (2019)) to estimate importance weights on the client side. For the privacy span $s$ in the user input $I$, the importance weight $w_{s}$ is calculated as the average attention received by $s$:
$w_{s}$ $=$ $\frac{1}{n}\frac{1}{n_{h}}\sum_{t=1}^{n}\sum_{h=1}^{n_{h}}\text{Attn}_{h}(s,i_{t}),$ (4)
where $n$ is the number of tokens in the input, $n_{h}$ is the number of attention heads in the lightweight model, $i_{t}$ is the $t$-th token of $I$, and $\text{Attn}_{h}(s,i_{t})$ denotes the attention score of $i_{t}$ attending to the privacy span $s$. To simplify the problem, we first consider constructing the meta vector for a single head $h$. The meta vector $\mathcal{R}_{h}$ on head $h$ is obtained by computing the weighted sum of the restoration vector $r_{s}^{h}$ of each privacy span $s$ on head $h$, normalizing the summation, and adding noise $\mathcal{N}$. The process is formulated as follows,
$Z_{h}$ $=$ $\frac{\sum_{s\in\mathcal{S}_{I}}w_{s}\cdot r_{s}^{h}}{||\sum_{s\in\mathcal{S}_{I}}w_{s}\cdot r_{s}^{h}||_{2}},$ (5)
$\mathcal{R}_{h}$ $=$ $Z_{h}+\mathcal{N},$ (6)
@@
where $\mathcal{S}_{I}$ denotes the set of privacy spans in the input $I$, $|\mathcal{S}_{I}|$ denotes the number of privacy spans and $Z_{h}$ represents the normalization of the weighted sum on head $h$. The injected noise $\mathcal{N}$ is sampling from the distribution $p(\mathcal{N})\propto\exp(-\epsilon\|\mathcal{N}\|)$, according to Feyisetan et al. (2020), where $\epsilon$ is the privacy hyperparameter.
To construct the meta vector for multiple heads, we first concatenate the restoration vectors on multiple heads. Then, we apply the weighted summary, normalization, and noise addition to the concatenated vector, as we do for a single edited head. After construction, the meta vector, along with the input with the privacy spans removed, are transmitted to the server for inference.
@@
We utilizes the meta vector to restore the missing privacy spans during inference on the input with the privacy spans removed, as illustrated in the lower right part of Figure 1. This operation is conducted on the server side.
Following activation steering methods (Li et al., 2023c; Chen et al., 2024), we apply the meta vector to the outputs of the edited attention heads to restore the privacy spans. Let ${\textbf{u}}_{h}$ represent the hidden state of the last token on head $h$ given the input with privacy spans removed, and $\mathcal{R}_{h}$ be the meta vector for head $h$, the hidden state of the last token on head $h$ after restoration $\bar{\textbf{u}}_{h}$ is denoted as:
$\bar{\textbf{u}}_{h}={\textbf{u}}_{h}+||\textbf{u}_{h}||_{2}\cdot\mathcal{R}_{h},\ \forall h\in\mathcal{H}_{c}.$ (7)
During inference, if a head belongs to the common top-K heads set $\mathcal{H}_{c}$, its hidden state should be modified using Eq 7.
5 Analysis of Privacy Budget
In this section, we analyze the privacy budget of DP variants and PrivacyRestore.
@@ 5 Analysis of Privacy Budget
The DP variants including CDP, LDP and $d_{\chi}$-privacy are constrained by a privacy budget that grows linearly with the length of the protected text.
The detail proof of Theorem 5.1 is presented in Appendix E. As the length of the protected text increases, the growing privacy budget makes these DP variants more vulnerable to adversarial attacks, thus compromising their robustness. We also provide empirical evidence demonstrating the linear growth problem of $d_{\chi}$-privacy in Section 6.3. We implement two types of attack, i.e., prompt injection attack (Perez and Ribeiro, 2022; Suo, 2024) and attribute inference attack (Li et al., 2022), across three privacy-preserving datasets. As shown in Figure 3(a) and 3(b), attack performance increases with the length of the protected text, highlighting the linear growth problem of the privacy budget in $d_{\chi}$-privacy.
Theorem 5.2.
PrivacyRestore fulfills $d_{\chi}$-privacy and provides a privacy budget of $\epsilon||Z^{\prime}-Z||$, where $\epsilon$ denotes privacy hyperparameter, and $Z^{\prime}$ and $Z$ represent any pair of normalized weighted sums of restoration vectors concatenated across all edited heads. The privacy budget of PrivacyRestore is independent of the length of protected text.
The detail proof of Theorem 5.2 is provided in Appendix F. The $\epsilon$ is privacy hyperparameter, independent of the length of protected text. The $||Z^{\prime}-Z||$ represents the distance between two vectors, also independent of the length of protected text. Since the privacy budget in PrivacyRestore is independent of the length of the protected text, our method effectively protects privacy even with longer protected text, inherently addressing the linear growth issue in DP variants. We also provide empirical evidence to support the theorem. We present the attack performance of PrivacyRestore across varying protected text lengths, as detailed in Section 6.3.
@@ 6.1 Experiments Setup
We evaluate our method in medical and legal domains. However, existing benchmarks, such as DDXPlus (Tchango et al., 2022) and NLICE (Al-Ars et al., 2023) for medical diagnosis, and SLJA (Deng et al., 2023b) for legal judgement, do not specify privacy spans in the input. To address this gap, we leveraged GPT-3.5 (Ouyang et al., 2022) to classify symptoms in DDXPlus/NLICE and case details in SLJA into sensitive and non-sensitive categories, treating the sensitive data as privacy spans. The classification prompt is shown in Appendix O.1. Based on classification results, we curated three privacy-preserving datasets: Pri-DDXPlus, Pri-NLICE and Pri-SLJA, with 149, 64 and 142 types of privacy spans, respectively. The process of dataset construction and statistical information can be found in Appendix B.
@@
The evaluation access both performance and inference efficiency. For performance evaluation, we employ MC1 and MC2 (Lin et al., 2021) to measure the model’s accuracy in selecting the correct answer among 4 options. Each sample in Pri-DDXPlus, Pri-NLICE, and Pri-SLJA is assigned with 4 options, including one correct and three incorrect options. The detailed calculation of MC1 and MC2 is outlined in Appendix J. We also evaluate the model’s generation ability using ROUGE-L and LLM-Judge (LLM-J)(Zheng et al., 2023a). For ROUGE-L, the reference text is the initial output produced by the backbone LLM. As ROUGE-L primarily focuses on n-gram overlap between generated text and reference texts, which may not fully capture the semantic meaning or overall quality of the generated content, we further use a LLM (i.e., GPT-3.5) to assess the quality of outputs considering relevance, clarity, and accuracy. The assessment prompt is shown in Appendix O.3. The LLM-J score ranges from 1 to 10, with higher scores indicating better quality. For inference efficiency, we use Throughput (TP), defined as the number of tokens generated per second, to evaluate the inference efficiency.
@@
To demonstrate the effectiveness of our method, we compare our model with following baselines. ${d_{\chi}}$-privacy. The client applies $d_{\chi}$-privacy (Feyisetan et al., 2020) mechanism on the entire input, which injects noise into tokens’ embedding and find the nearest tokens to replace the initial tokens. ${d_{\chi}}$-privacy on privacy spans. The client employs $d_{\chi}$-privacy (Feyisetan et al., 2020) mechanism on privacy spans in the input, rather than the entire input. Paraphrase. According to Mattern et al. (2022b); Utpala et al. (2023), clients can use generative models to paraphrase original inputs, achieving effects similar to DP.
Implementation Details.
We use Llama2-chat-7b (Touvron et al., 2023) as the LLM backbone on the server side, and BERT-base (Devlin et al., 2019) on the client side for weight estimation, as described in Section 4.2. For fair comparison, we utilize flan-t5-base model (Chung et al., 2024) on the client side for paraphrasing in the Paraphrase baseline as its model size is comparable to that of BERT-base. During restoration vector training, the LLM parameters remain fixed, and we train the restoration vectors for 5 epochs with a batch size of 1. The optimal number of edited heads $K$ is 175 for Pri-DDXPlus/Pri-SLJA and 125 for Pri-NLICE. The search process is shown in Section I. To evaluate the generation capabilities, we utilize GPT-3.5 to assess the generated outputs. The prompts are detailed in Appendix O.2. To evaluate inference efficiency, we use the greedy search decoding strategy and set the max generation length to 256 during generation.
@@
The hyperparameters related to privacy protection strength are $\epsilon$ for $d_{\chi}$-privacy (on privacy spans) and PrivacyRestore, and $\tau$ for paraphrase. For fair comparison, we ensure all methods under the same privacy budget. We show the calculation process of determining values of $\epsilon$ and $\tau$ for different methods on different datasets in Appendix H. The values of $\epsilon$ and $\tau$ for different privacy-preserving methods are shown in Table 5 in the Appendix.
@@ 6.2 Main Results
Pri-DDXPlus $d_{\chi}$-privacy 28.79$\pm$0.02 30.26$\pm$0.01 17.97$\pm$0.00 1.17$\pm$0.00 37.45$\pm$0.01
$d_{\chi}$-privacy on privacy spans 44.71$\pm$0.29 42.36$\pm$0.00 29.17$\pm$0.04 3.31$\pm$0.00 33.21$\pm$0.00
Paraphrase 27.92$\pm$0.56 28.56$\pm$0.07 18.04$\pm$0.01 1.23$\pm$0.00 35.42$\pm$0.67
PrivacyRestore 62.97$\pm$0.00 60.19$\pm$0.00 27.24$\pm$0.26 4.47$\pm$0.00 26.09$\pm$0.08
Pri-NLICE $d_{\chi}$-privacy 29.08$\pm$0.00 29.72$\pm$0.00 15.68$\pm$0.02 1.41$\pm$0.00 38.30$\pm$0.00
$d_{\chi}$-privacy on privacy spans 30.00$\pm$0.09 31.46$\pm$0.00 22.97$\pm$0.00 3.01$\pm$0.00 35.73$\pm$0.57
Paraphrase 28.46$\pm$0.02 29.15$\pm$0.03 16.15$\pm$0.01 1.62$\pm$0.00 37.22$\pm$0.07
PrivacyRestore 62.23$\pm$1.70 57.94$\pm$0.09 24.42$\pm$0.81 3.67$\pm$0.01 32.33$\pm$0.01
Pri-SLJA $d_{\chi}$-privacy 16.66$\pm$0.37 17.57$\pm$0.04 23.35$\pm$0.00 2.08$\pm$0.00 36.83$\pm$0.03
$d_{\chi}$-privacy on privacy spans 24.23$\pm$1.69 26.63$\pm$0.67 40.10$\pm$0.00 4.54$\pm$0.00 36.16$\pm$0.00
Paraphrase 16.21$\pm$0.02 17.52$\pm$0.02 24.90$\pm$0.01 2.07$\pm$0.01 31.31$\pm$0.05
PrivacyRestore 35.47$\pm$1.48 35.41$\pm$0.64 37.56$\pm$0.06 5.25$\pm$0.00 30.73$\pm$0.04
@@
PrivacyRestore achieves best scores in MC1/2 and LLM-J compared to other privacy-preserving methods. In terms of the ROUGE-L evaluation metric, PrivacyRestore achieve the best result in Pri-NLICE while ranking second in the other two datasets. This discrepancy likely stems from ROUGE-L’s dependence on n-gram overlap between the reference text and the generated output, which does not fully reflect the quality of generated outputs. As demonstrated by the examples in Figure 6 and Appendix N, PrivacyRestore often generates outputs with different sentence structures while still providing accurate answers. Consequently, our method achieves slightly lower ROUGE-L scores but significantly higher LLM-J scores compared to $d_{\chi}$-privacy on privacy spans. Furthermore, the ROUGE-L metric displays larger variance than the LLM-J metric, potentially due to its sensitivity to expression rather than the underlying meaning of the generated output.
As for inference efficiency, $d_{\chi}$-privacy achieves the highest throughput. In contrast, $d_{\chi}$-privacy on privacy spans requires prior identification of privacy spans, while paraphrase necessitates rephrasing the user input on the client side, leading to delays. PrivacyRestore also requires additional time for the prior identification and removal of privacy spans, along with constructing the meta vector on the client side. However, its throughput can reach nearly 70% on Pri-DDXPlus and 80% on Pri-NLICE and Pri-SLJA, relative to the best results.
@@ 6.3 Empirical Privacy Protection Results
We not only provide a theoretical privacy proof of our method in Section 5, but also implement two attack methods to empirically evaluate the privacy protection capability of our approach. PrivacyRestore sends only the meta vector and the incomplete input with privacy spans removed. It is less likely for adversaries to infer privacy spans from incomplete input. Adversaries can only attack by intercepting the meta vector and inferring the corresponding privacy spans from it. Therefore, we implement the embedding inverse attack (Li et al., 2023b; Morris et al., 2023) and attribute inference attack (Li et al., 2022), both commonly used methods for attacking embeddings. For $d_{\chi}$-privacy, $d_{\chi}$-privacy and paraphrase methods, we obtain the hidden state of the last token in the last layer in Llama2-chat-7b as the embeddings. The embedding inverse attack utilizes ROUGE-L as the evaluation metric, while the attribute inference attack employs F1 to assess attack performance. Detailed description about both attack methods and evaluation metrics can be found in Appendix M.
@@
Figure 2: Results of all methods under embedding inverse attack and attribute inference attack under different privacy hyperparameters $\epsilon$ on three datasets.
Different Privacy Hyperparameter $\epsilon$.
We evaluate the attack performance of the embedding inverse attack and attribute inference attack for PrivacyRestore and other baselines. Additionally, we present the attack results without any privacy protection, which serves as the upper bound of attack performance. The values of $\epsilon$ on x-axis in Figure 2 represent the values used in PrivacyRestore. Equivalent values of $\epsilon$ and $\tau$ for other baselines are provided in the Appendix H to ensure the same privacy budget. As shown in Figure 2, although the privacy budget is the same, PrivacyRestore demonstrates better privacy protection performance, evidenced by its lower ROUGE-L and F1 scores. All privacy-preserving methods effectively protect privacy compared to the upper bound of no protection.

### v5

@@ 4.2 Inference Stage
According to the principle of “Information Self-Determination Right”, users should identify the privacy spans in their input by themselves, because the definition of privacy varies from person to person. For each privacy span, the client employs a lightweight model (e.g. BERT Devlin et al. (2019)) to classify it into a specific type within the predefined privacy span type set $\mathcal{C}$. For example, the privacy span “My stools sometimes is black” will be classified into the predefined privacy span type black stools. Due to the long-tailed distribution of privacy spans, our predefined type set can cover the majority of privacy spans. Even when encountering privacy spans of out-of-set types, classifying these rare spans into the types of the predefined set can still be effective, as shown in §6.6.
Then the client should aggregate those restoration vectors corresponding to the privacy spans into a single meta vector. Transmitting a singular meta vector enhances privacy protection compared to multiple vectors, as it prevents potential information leakage regarding the quantity of privacy-sensitive segments.
@@
However, Equal Weighted Aggregation (EWA) may weaken the influence of critical privacy spans and amplify the effect of irrelevant ones. Therefore, we propose a novel method called Attention-aware Weighted Aggregation (AWA) which estimates a weight for each privacy span and then takes the weighted sum of corresponding restoration vectors as the meta vector. Given computational constraints on client devices, we employ a lightweight BERT model to evaluate privacy span significance by calculating the mean attention score $w_{s}$ across all attention heads and tokens within the input query. This metric quantifies the relative importance of each privacy-related span.
Considering that each privacy span type $c$ will have multiple restoration vectors $r_{h}^{c}$ across all edited heads in $\mathcal{H}_{k}$, we first concatenate these restoration vectors from multiple heads to form $r^{c}$ for privacy span type $c$. Then, we compute the meta vector $\mathcal{R}$ by calculating the weighted sum of the restoration vector $r^{c}$, normalizing the summary, and adding noise $\mathcal{N}$ for privacy protection. The process is formulated as follows:
$r^{c}$ $=$ $\text{Concat}(r_{1}^{c},r_{2}^{c},...,r_{h}^{c}),$ (1)
$Z$ $=$ $\frac{\sum_{s\in\mathcal{S}_{q}}w_{s}\cdot r^{c}}{||\sum_{s\in\mathcal{S}_{q}}w_{s}\cdot r^{c}||_{2}},$ (2)
$\mathcal{R}$ $=$ $Z+\mathcal{N},$ (3)
@@
where $s$ represents the privacy span of type $c$, $\mathcal{S}_{q}$ denotes all privacy spans in the user query $q$ and $Z$ represents the normalization of the weighted sum, which can also be viewed as the meta vector without protection. The injected noise $\mathcal{N}$ is sampling from the distribution $p(\mathcal{N})\propto\exp(-\epsilon\|\mathcal{N}\|)$, to achieve the $d_{\chi}$-privacy mechanism, where $\epsilon$ is the privacy hyperparameter (Feyisetan et al., 2020).
After construction, the meta vector $\mathcal{R}$ and the incomplete query $\hat{q}$ (with privacy spans removed) are transmitted to the server for inference.
@@
We utilize the meta vector $\mathcal{R}$ to restore the information in the removed privacy spans during inference, as illustrated in the lower right part of Figure 1. This operation is conducted on the server side.
Following activation steering methods (Li et al., 2023c; Chen et al., 2024), we apply the meta vector to the outputs of the edited attention heads to achieve restoration. Let ${\textbf{u}}_{h}$ represent the hidden state of the last token on head $h$ given the incomplete user query $\hat{q}$ and $\mathcal{R}_{h}$ denotes a part of the meta vector $\mathcal{R}$ for head $h$. Then the hidden state of the last token on head $h$ after restoration, denoted as $\bar{\textbf{u}}_{h}$, can be computed by:
$\bar{\textbf{u}}_{h}={\textbf{u}}_{h}+||\textbf{u}_{h}||_{2}\cdot\mathcal{R}_{h},\ \forall h\in\mathcal{H}_{k}.$ (4)
During inference, if a head belongs to the common top-K heads set $\mathcal{H}_{k}$, its hidden state should be modified using Eq 4. To prevent privacy leakage from the generated output, we employ sampling-based generation, which is protected by the Exponential Mechanism (Utpala et al., 2023a).
5 Privacy Guarantee Analysis
In this section, we analyze the privacy guarantees and privacy budget of PrivacyRestore.
Our approach transmits only a privacy-free incomplete query and a meta vector secured by $d_{\chi}$-privacy mechanism. Therefore, even if attackers steal both the incomplete user query and the meta vector during transmission or even hack the server, they still cannot infer any user privacy. Furthermore, the confidentiality of the server’s LLM parameters and edited head set $\mathcal{H}_{k}$ prevents attackers from reconstructing the generation process using intercepted elements, ensuring robust security against privacy breaches. Then we analyze the privacy budget of our method, as follows:
@@ 5 Privacy Guarantee Analysis
PrivacyRestore fulfills $d_{\chi}$-privacy and provides a privacy budget of $2\epsilon$, where $\epsilon$ denotes privacy hyperparameter. The privacy budget of PrivacyRestore is independent of the length of the protected text.
Pointed by Mattern et al. (2022b), directly applying $d_{\chi}$-privacy mechanism to all tokens in the user query, for privacy protection, suffers from the linear growth problem of privacy budget. In contrast, our method ensures that the privacy budget remains constant at $2\epsilon$, independent of the length of protected text. We also provide empirical evidence to demonstrate that our approach effectively addresses the linear growth problem of the privacy budget encountered in $d_{\chi}$-privacy in Section 6.4. Detailed proof of Theorem 5.1 is provided in Appendix H.
@@ 6.1 Experiments Setup
To evaluate our method, we construct three privacy-preserving datasets covering the medical and legal domains: Pri-DDXPlus, Pri-NLICE, and Pri-SLJA. The detailed process of dataset construction and statistical information can be found in Appendix J.
@@
The evaluation assesses both performance and inference efficiency. For performance evaluation, we use MC1/MC2 (Zhang et al., 2024), ROUGE-L (Lin, 2004), and LLM-Judge (LLM-J) (Zheng et al., 2023a) metrics. For inference efficiency, we use the Throughput (TP) metric. The details of these metrics and their corresponding calculation processes are provided in Appendix L.1.
@@
To demonstrate the effectiveness of our method, we compare our model with the following baselines: No Protection, No Restoration, ${d_{\chi}}$-privacy (Feyisetan et al., 2020), ${d_{\chi}}$-privacy on privacy spans and Paraphrase Mattern et al. (2022b); Utpala et al. (2023b). A detailed introduction to these baseline methods is provided in Appendix L.2.
@@
The hyperparameters related to privacy protection strength are $\epsilon$ for $d_{\chi}$-privacy (on privacy spans) and PrivacyRestore, and $\tau$ for paraphrase. For a fair comparison, we ensure all methods are under the same privacy budget. We show the calculation process of determining values of $\epsilon$ and $\tau$ for different methods on different datasets in Appendix M.
@@ 6.2 Main Results
Pri-DDXPlus No Restoration (lower bound) 33.57±0.00 32.49±0.01 25.19±0.43 3.21±0.01 40.86±0.01
No Protection (upper bound) 64.88±0.01 61.48±0.03 100.00±0.00 5.58±0.03 41.08±0.09
$d_{\chi}$-privacy 28.79±0.02 30.26±0.01 17.97±0.00 1.17±0.00 37.45±0.01
$d_{\chi}$-privacy on privacy spans 44.71±0.29 42.36±0.00 29.17±0.04 3.31±0.00 33.21±0.00
Paraphrase 27.92±0.56 28.56±0.07 18.04±0.01 1.23±0.00 35.42±0.67
PrivacyRestore 62.97±0.00 60.19±0.00 27.24±0.26 4.47±0.00 26.09±0.08
Pri-NLICE No Restoration (lower bound) 27.07±1.98 28.63±2.23 16.90±0.51 1.61±0.03 41.08±0.01
No Protection (upper bound) 80.30±0.38 77.60±1.23 100.00±0.00 5.90±0.04 41.44±0.04
$d_{\chi}$-privacy 29.08±0.00 29.72±0.00 15.68±0.02 1.41±0.00 38.30±0.00
$d_{\chi}$-privacy on privacy spans 30.00±0.09 31.46±0.00 22.97±0.00 3.01±0.00 35.73±0.57
Paraphrase 28.46±0.02 29.15±0.03 16.15±0.01 1.62±0.00 37.22±0.07
PrivacyRestore 62.23±1.70 57.94±0.09 24.42±0.81 3.67±0.01 32.33±0.01
Pri-SLJA No Restoration (lower bound) 24.92±0.98 25.97±1.12 31.02±0.16 4.43±0.01 39.14±0.09
No Protection (upper bound) 69.57±0.61 67.58±0.43 100.00±0.00 5.44±0.03 39.49±0.13
$d_{\chi}$-privacy 16.66±0.37 17.57±0.04 23.35±0.00 2.08±0.00 36.83±0.03
$d_{\chi}$-privacy on privacy spans 24.23±1.69 26.63±0.67 40.10±0.00 4.54±0.00 36.16±0.00
Paraphrase 16.21±0.02 17.52±0.02 24.90±0.01 2.07±0.01 31.31±0.05
PrivacyRestore 35.47±1.48 35.41±0.64 37.56±0.06 5.25±0.00 30.73±0.04
@@
PrivacyRestore achieves best scores in MC1/2 and LLM-J compared to other privacy-preserving methods. In terms of the ROUGE-L evaluation metric, PrivacyRestore achieve the best result in Pri-NLICE while ranking second in the other two datasets. This discrepancy likely stems from ROUGE-L’s dependence on $n$-gram overlap between the reference text and the generated output, which does not fully reflect the quality of generated outputs. As demonstrated by the examples in Figure 8 and Appendix W, PrivacyRestore often generates outputs with different sentence structures while still providing accurate answers. Consequently, our method achieves slightly lower ROUGE-L scores but significantly higher LLM-J scores compared to $d_{\chi}$-privacy on privacy spans. Furthermore, the ROUGE-L metric displays larger variance than the LLM-J metric, potentially due to its sensitivity to expression rather than the underlying meaning of the generated output. Shown in Table 1, No Protection servers as the performance upper bound for all privacy-preserving methods while No Restoration servers as the performance lower bound. Our method significantly outperforms No Restoration and is even comparable to No Protection, strongly validating the effectiveness of our approach.
Although PrivacyRestore incurs slight latency from client-side privacy span identification and meta-vector construction, its throughtput attain nearly 70% of the best results, which is acceptable. We provide further analysis and additional experimental results in Appendix V.
@@ 6.3 Empirical Privacy Protection Results
In this section, we implement attack methods to empirically show that our approach offers superior privacy protection compared to baselines, both for user inputs and model outputs.
Privacy Protection Evaluation on Inputs.
In this section, we implement the embedding inverse attack (Li et al., 2023b; Morris et al., 2023) and attribute inference attack (Li et al., 2022) to attack the inputs of our method and other baselines, including the meta vector and the privacy-free incomplete user query. As shown in Figure 2, as the privacy budget increases, the privacy protection capability of all privacy-preserving methods decreases. However, PrivacyRestore consistently outperforms others across all privacy budgets, as indicated by its lower ROUGE-L and F1 scores.
@@
Figure 2: Results of embedding inverse attack and attribute inference attack for all baselines under different privacy hyperparameters $\epsilon$ on Pri-DDXPlus.
Privacy Protection Evaluation on Outputs.
We use the sampling-based method to generate the outputs on the server. As demonstrated by Utpala et al. (2023a); Mattern et al. (2022b), sampling-based generation satisfies the Exponential Mechanism (McSherry and Talwar, 2007), which can effectively prevent privacy leakage from the generated outputs. We implement the embedding inversion and attribute inference attacks for the generated outputs under various generation temperatures and also count the occurrence of privacy spans in the outputs. As shown in Table 2, the attack performance remains consistently low, demonstrating that sampling-based generation effectively prevents privacy leakage from the generated outputs. Implementation details of these attack methods can be found in Appendix O and Appendix Q.
We also provide a biref theoretical proof of our method’s output protection in Appendix I.
Temperature 0.75 1.0 1.25 1.5 1.75
EIA(ROUGE-L) 0.037 0.038 0.035 0.035 0.037
AIA(F1) 0.096 0.097 0.092 0.092 0.097
Occurrence 0.031 0.030 0.030 0.029 0.031
Table 2: Analysis of output privacy leakage from outputs on Pri-DDXPlus dataset. EIA denotes embedding inverse attack. AIA indicates attribute inference attack. Occurrence metric directly counts the frequency of privacy spans in the generated output. We primarily use a temperature of 1.0 during generation in the other experiments.
6.4 Privacy Protection for Long Text
In this section, we implement attack methods for both the $d_{\chi}$-privacy baseline and our PrivacyRestore approach under varying protected text lengths, illustrating robust privacy and addressing the linear growth of the privacy budget in $d_{\chi}$-privacy.
For $d_{\chi}$-privacy, we randomly select a proportion of tokens as $d_{\chi}$-privacy Percentage to protect—higher percentages yield longer protected text. Shown in Figures 3(a) and 3(b), both prompt injection and attribute inference attacks are implemented with attack performance increases with the percentage. It is caused by the linear growth problem of the privacy budget encountered in $d_{\chi}$-privacy, as raised by Mattern et al. (2022b).
For PrivacyRestore, a proportion of privacy spans is selected for protection, defined by the Privacy Span Ratio $\alpha$, with larger $\alpha$ indicating more spans. Shown in Figures 3(c) and 3(d), aside from the embedding inverse attack on the Pri-NLICE dataset, attack performance remains stable across different $\alpha$ values. These results confirm that our method provides strong privacy protection, even as text length increases.